OpenStudy (anonymous):

5x+2/2=2x-1 (slash is a fraction)

OpenStudy (anonymous):

\[\frac{ 5x+2 }{ 2 }=2x-1\] Here it is @GhostOfArtyom @Pocarii @Ashleyisakitty @bohotness @bibby @Bossimbacon Help pls

OpenStudy (anonymous):

Multiply the 2 both sides

OpenStudy (anonymous):

10x+ 4= 4x-2

OpenStudy (anonymous):

Subtract 4x on both sides

OpenStudy (anonymous):

6x + 4 = -2

OpenStudy (anonymous):

Subtract 4 to both sides

OpenStudy (anonymous):

6x= -6 Divide by 6 X= - 1

OpenStudy (anonymous):

Dang I could have helped, sorry I wasnt looking.

OpenStudy (anonymous):

@GhostOfArtyom Its fine...

OpenStudy (anonymous):

@artfan96 thanks...